OpenStudy (anonymous):

What is the sum of the interior angles of a hexagon? 1,260° 1,080° 900° 720°

OpenStudy (mertsj):

how many triangles can you divide a hexagon into?

OpenStudy (mertsj):



hero (hero):

360/6 = 60 180 - 60 = 120 120*6 = 720

hero (hero):

1. Take 360 degrees in a circle, divide by the number of angles on the interior of a polygon, (in this casesix). This subsequently gives you the measure of one of the exterior angles of a hexagon. 2. Since the exterior and interior angles of of a polygon are supplementary, take 180 degrees subtract the measure of the exterior angle to get the measure of the interior angle. 3. Multiply the measure of the interior angle by the number of sides in the polygon. This subsequently gives you the sum of the interior angles of a polygon.

hero (hero):